Use two equations in two variables to solve the application. See Example 1. (Objective 1)
Chayla and Lena pool their resources to buy several lottery tickets. They win $370,000! They agree that Lena should get $70,000 more than Chayla, because she gave most of the money. How much will Chayla

Answer by ikleyn(52786)

.

    L + C = 370000    (1)

    L - C =  70000    (2)


Add the equations

    2L = 440000

     L = 440000/2 = 220000.


It is how much Lena gets.


The rest  370000 - 220000 = 150000 dollars go to Chayla.

Solved.

Again, I can not see Example 1. (Objective 1) and do not understand for what reason this instruction is included to the post.